OpenStudy (anonymous):

The half-life of a specific radioactive compound is 15 years. The equation used to model this half-life is y = 200(1/2)t/15, where t represents time. Currently, there are 200 pounds of this substance remaining. Determine approximately how much of the substance remains after 82 years. Round to the nearest tenth of a year.

OpenStudy (anonymous):

replace \(t\) by \(82\) and use a calculator

OpenStudy (anonymous):

that is, compute \[\large 200\times \left(\frac{1}{2}\right)^{\frac{82}{15}}\]

OpenStudy (anonymous):

i got 546.6666667

OpenStudy (anonymous):

no it should be small

OpenStudy (anonymous):

make sure to raise one half to the power first, then multiply by 200 last

OpenStudy (anonymous):

so its 5

OpenStudy (anonymous):

says round to the nearest tenth